1. You have the following information for a firm: 2015 2016 2017 2018 Competitor 19163 21316 22387 24720 15790 Sales Net Income Dividends per share 0.68 0.85 1.05 1.32 568 0.43 2757 2818 2885 4518 1,513 1,487 1,462 1,394 245.8 14.39% 13.22% 12.89% 18.28% 3.60% 1.15 Shares Net Margin Asset Turnover Financial Leverage 2.13 2.43 2.63 20.54 4.51 Return on Equity 47.4% 47.8% 52.9% 384.3% 18.7% EPS 1.54 1.49 .56 1.02 1.82 1.90 1.97 3.24 2.24 a. For the most recent year, what part of EPS is paid out as dividends? C- b. Review the Du Pont identity for this firm - what is driving the ROE? How do they compare to the competitor? c. What is the average change in sales for this firm?
Answer #1

a)Most recent year here means 2018
dividends paid out of EPS= Dividend per share/EPS
=1.32/3.24=40.74%

b)ROE= Net income/equity
=(net iccome/sales)*(sales/assets)*(assets/equity)
=net margin*asset turnover*financial leverage
for recent year 2018 te ROE of firm is
=18.28%*1.02*20.54=384.3%
for the competitor it is
=3.6%*1.15*4.51=18.7%
we can see for firm it is very highe ROE because of high net margin and the financial leverage it carries

c)we will see the change in sales each year from 2015 to 2018 and do average of this
change in sales=(final/initial)-1
15 to 16=(21316/19163)-1=11.24%
16 to 17=(22387/21316)-1=5.02%
17 to 18=(24720/22387)-1=10.42%
Average change=(11.24%+5.02%+10.42%)/3=8.89%